please answer its urgent.

develop f(z)=(z(z-3)) into a laurent serkes valid for the following domains

7) 0 < 1 2 -3/ <3 6) 1८11-4/<4
develop g(z)= 1/((z-1)(z-2)) into a laurent series valid for the following domains

9) 0시레시 10) 0<l2-2시
develop h(z)= z/((z+1)(z-2)) into a laurent series valid for the following domains
) ۵ < ( 2 + ( ( 3 (2) 02 ( 2 -2) 3.
Answer #1

9 १(-) = 7(2-3) 7 domain ०/-3/ 23 f(z) is every where analytic Laurent series expension < (7-3) 22-37 8) domain ।</z-4/२५ herZ-) 자2 ID) 0시2221시 (c) 굶 224 파리시 긽 (+19) ㅋ 2 - [1- (Z-2) + (z-2) -l2-294 - - ] 故 깨다29에 is the lawrent series of 912) in 0</2-h녀 래 = 2 +2 음 3 0 221~3 0< 시 해 +를 ㅢ ㅎ 첨가3 +3t를 긽 , t를 를 싫 1 시 내 끓릉긁 - + - --] + 글 층한 t5 (-1)() m 0시2213 - 2 M20